The question is: Write the equation of the line in standard form through the point (5,-1) with the slope of 3/4. So far what I have done is substituted the point into the standard form equation and it came out like this: 5a - 1b=c i don't know what see is and what further work i should do to get the answer.
Answer by stanbon(75887)

The question is: Write the equation of the line in standard form through the point (5,-1) with the slope of 3/4.
--------------------
You know that m = 3/4 and that y=-1 when x=5
---
Substitute into y = mx + b to solve for "b":
-1 = (3/4)5 + b
(-4/4) = (15/4) + b
b = -19/4
===========
Equation:
y = (3/4)x - 19/4
==================
Cheers,
Stan H.